OpenStudy (anonymous):

Rewrite the following in simplest radical form. √27w^8

OpenStudy (lalaly):

\[=\sqrt{27w^8}=(27w^8)^{\frac{1}{2}}=(3^3w^8)^\frac{1}{2}=3w^4 \sqrt3\]

OpenStudy (anonymous):

Let g and f be two functions defined by g(x)=5-4x^2 f(x)=9-5x for any real number . Find g/f(-5) and then find all values that are NOT in the domain of g/f . If there is more than one value, separate them with commas.

OpenStudy (lalaly):

what is this?